# Permutation invariant matrix quantum thermodynamics

### v1

@@ 10 Summary and Outlook
Computations of the degeneracies at low energies, with the help of representation theoretic formulae for the micro-canonical data, gives evidence that the features of negative specific heat capacity and ensemble inequivalence also occur in certain multi-matrix and tensor systems. Hence such systems have a low temperature region with negative micro-canonical specific heat and a high temperature positive specific heat. In particular we found that the charge zero sector of the Hermitian two-matrix model (with the two matrices rotating under an additional $U(1)$ or $SO(2)$) again has a Hagedorn type transition at $T_{H}=\frac{1}{\ln 2}$ but the low temperature approach to this transition is described in eqn (6.48) in contrast to eqn (8.5) for the two matrix model.
@@ Appendix D SAGE computation for zero charge sector of 2-matrix models
Appendix E SAGE computation for negative SHC in multi-matrix model with $d=n$ scaling
@@
# representation associated with Young diagram p having n boxes
@@
def DimmGenNum  (d , p, n  ) :
    return prod ( prod ( ( d - i +  k  ) for k in range (p[i])  ) for i in range ( len ( p )))
def DimmGen  (d , p, n  ) :
    return DimmGenNum  (d , p, n  )*dimension(p)/factorial (n)
@@
## Counting states at energy n in d-matrix harmonic oscillator where d is set to n, using character sum
@@
def Zmultmatdeqn (  n , N ) :
    L = len ( Partitions(n).list() )
@@
                 if len( Partitions(n).list()[i]) < N+1  :
                        S = S + (s(P[i]).scalar(pp(P[j])))^2/P[j].centralizer_size()*n^(len (P[j]))

### v2

@@ 10 Summary and Outlook
Computations of the degeneracies at low energies, with the help of representation theoretic formulae for the micro-canonical data, gives evidence that the features of negative specific heat capacity and ensemble inequivalence also occur in certain multi-matrix and tensor systems with continuous gauge symmetries. Hence such systems have a low temperature region with negative micro-canonical specific heat and a high temperature positive specific heat. In particular we found that the charge zero sector of the Hermitian two-matrix model (with the two matrices rotating under an additional $U(1)$ or $SO(2)$) again has a Hagedorn type transition at $T_{H}=\frac{1}{\ln 2}$ but the low temperature approach to this transition is described in eqn (6.48) in contrast to eqn (8.5) for the two matrix model.
@@ Appendix D SAGE computation for zero charge sector of 2-matrix models
Appendix E SAGE computation for negative SHC in multi-matrix model with $d=k$ scaling
@@
# representation associated with Young diagram p having k boxes
@@
def DimmGenNum  (d , p,  k  ) :
    return prod ( prod ( ( d - i +  l  ) for l in range (p[i])  ) for i in range ( len ( p )))
def DimmGen  (d , p, k  ) :
    return DimmGenNum  (d , p, k  )*dimension(p)/factorial (k)
@@
## Counting states at energy k in d-matrix harmonic oscillator where d is set to k, using character sum
@@
def Zmultmatdeqn (  k , N ) :
    L = len ( Partitions(k).list() )
@@
                 if len( Partitions(k).list()[i]) < N+1  :
                        S = S + (s(P[i]).scalar(pp(P[j])))^2/P[j].centralizer_size()*k^(len (P[j]))